Calculate the approximate pH of a 1% w/v Ammonium Chloride Solution. MW=53.5 pkb of Ammonium Hydroxide= 4.76
Answer #1

concentration = 1 % (w/v)

1 %(w/v) = 1 x 10 / 53.5

               = 0.187 M

pKb = 4.76

pH = 7 - 1/2 (pKb + log C)

      = 7 - 1/2 (4.76 + log 0.187)

pH = 4.98